OpenStudy (anonymous):

Use a system of equations to solve the following problem: Douglas High School has a total of 89 boys and girls who play sports. If the number of boys is 8 more than twice the number of girls, how many boys play sports at this high school? 54 27 81 62

OpenStudy (anonymous):

Well first. B+G=89 2G+8=B

OpenStudy (anonymous):

i think it is 62 am i correct?

OpenStudy (raden):

B + G = 89 ... (1) B = 2G + 8 ... (2) subtitute (2) to (1), giving us B+G=89 2G+G+8=89 3G + 8 = 89 subtrac by 8 from both sides, 3G + 8-8 = 89-9 3G = 81 divide by 3 from both sides, 3G/3 = 81/3 G = 27 ... (3) subtitute (3) to (2) B = 2G + 8 B = 2(27) + 8 B = 54 + 8 B = 62